hi can you pls help me to my assignment?

2x+y=1
i need you help

Answer by tutorcecilia(2152)

This is the equation of a line. To convert this to the slope-intercept format of the equation of a line, solve for y:
2x+y=1
2x-2x+y=1-2x
y=-2x+1 (switch terms)
This means that for any (x,y) points on this line, the slope will be -2 and the y-intercept will be 1.